Did I complete this right?
2/5=x-2/20
I started by multiplying 5 through the problem:
(5)(2/5)=(5)(x-2/20) and got: 2=5x-10/100!
Is this right or did I miss a step? Any advice would be great!
Thanks!
ac

Answer by stanbon(48579)

2/5=x-2/20
I started by multiplying 5 through the problem:
(5)(2/5)=(5)(x-2/20) and got: 2=5x-10/100!
You should have gotten:
2=5x-10/20 because the "5" multiplies the 2 but not the 20.
Then:
2=5x-(1/2)
Multiply thru by 2 to get:
4=10x-1
10x=5
x=1/2
Comment:
The Least Common Denominator of your original problem
was 20. You should have multiplied both sides by 20
to clear away all denominators, as follows:
20(2/5)=20(x-2/20)
8=20x-2
20x=10
x=1/2
Cheers,
Stan H.